OpenStudy (anonymous):

HELP ME ASAP!!! solve by graphing 5x^2+20x+15=0

OpenStudy (anonymous):

A

OpenStudy (anonymous):

B

OpenStudy (anonymous):

C

OpenStudy (anonymous):

D

OpenStudy (agent0smith):

since the number in front of x^2 in 5x^2+20x+15=0 is positive, the graph will open upwards... so you can eliminate the graph that opens down. You can also divide both sides of 5x^2+20x+15=0 by 5 to make it easier to solve.

OpenStudy (agent0smith):

(divide *everything* by 5)

OpenStudy (anonymous):

so wait were solving the equation by x right ?

OpenStudy (agent0smith):

Yes, solving it for x. But you can divide it by 5 first, then we'll solve it.

OpenStudy (anonymous):

so it would be x^2+20x+15=5 ?

OpenStudy (agent0smith):

You need to divide every term of 5x^2+20x+15=0 by 5. Every term, everything.

OpenStudy (agent0smith):

\[\Large \frac{ 5x^2+20x+15 }{ 5 }=\frac{ 0 }{5}\]

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

so its x^2+4x+3=5

OpenStudy (agent0smith):

Correct on the left side. What's 0/5? What's 0 divided by anything?

OpenStudy (agent0smith):

It's not 5...

OpenStudy (anonymous):

oh haha my bad

OpenStudy (anonymous):

okay so it would be x^2+4x+3=0 and then do you just + 3 on both sides ?

OpenStudy (agent0smith):

No leave everything where it is, now try to factor it.

OpenStudy (anonymous):

how would you do that ? :P

OpenStudy (agent0smith):

Find two numbers that multiply to the last number, 3, and add up to 4

OpenStudy (anonymous):

3*1 is 3 and 3+1 is 4

OpenStudy (agent0smith):

Yep now put them into your (x+3)(x+4) and so find which graph has zeroes -3 and -4

OpenStudy (anonymous):

so it would be D ?

OpenStudy (agent0smith):

Oh, yes sorry, (x+3)(x+1) not what i wrote above, you're correct about D :)

OpenStudy (anonymous):

okay thank you :)